A torque of 12 N m is applied to a solid, uniform puck of radius 0.50 m, causing the puck to accelerate at 5.7 rad/s2. What is the mass of the disk?


1
Expert's answer
2022-03-21T12:45:17-0400

Answer

Torque

"\tau=I\alpha"

So mass

"M=\frac{2\tau}{R^2\alpha}\\=\frac{2*12}{0.5*0.5*5.7}\\=16.84Kg"